OpenStudy (anonymous):

Combine like terms. 7/8x - 1/2x A. 1/8x B. 3/8 C. 3/8x D. x

OpenStudy (shamil98):

First, find a common denominator for both fractions, in this case we can use 8. So we leave 7/8 x unchanged, by multiply -1/2 x by 4/4 \[\huge \frac{ 7 }{ 8 } x - \frac{ 1 }{ 2 }x \] \[\Huge \frac{ 7 }{ 8 } x - \frac{ 4 }{ 8 }x \] Can you simplify from here?

OpenStudy (anonymous):

THNX!!! :D